Calculate the number of sides a regular polygon has if an interior angle is five times of an exterior angle?please show working
Answer by Alan3354(69443)

The interior + exterior = 180
I = 5E
E + I = 180
5E + E = 180
E = 36
=========
The sum of exterior angles for ALL polygons is 360 degs
360/36 = 10 sides